Company A, is an unlevered firm with expected annual earnings before taxes of $23,268,160 in perpetuity. The current required return on the firm’s equity is 15 percent, and the firm distributes all of its earnings as dividends at the end of each year. The company has 1.34 million shares of common stock outstanding and is subject to a corporate tax rate of 40 percent. The firm is planning a recapitalization under which it will issue $16,379,560 of perpetual 9.4 percent debt and use the proceeds to buy back shares. Use the APV method to calculate the company value after the recapitalization plan is announced.

74,825,504

97,760,000

67,360,000

67,360,000
Answer #1

Net income = Expected annual earnings before tax - tax @ 40%

= $23,268,160 - ($23,268,160 * 40%)

= $23,268,160 - $9,307,264

= $13,960,896

Value of company as unlevered firm = Net income/required return = $13,960,896/15% = $93,072,640

Interest on debt =  $16,379,560 * 9.4% = $1,539,678.64

Interest tax shield = $1,539,678.64 * (40%) = $615,871.50

PV of tax shield = $615,871.50/9.4% = $6,551,824.50

Company value after recapitalization = Value of company as unlevered firm + PV of tax shield

= $93,072,640 + $6,551,824.50

= $99,624,464.5

don't what wrong with options.